# CiFlow: Dataflow Analysis and Optimization of Key Switching for Homomorphic Encryption

### v2

@@ VI-C Sensitivity Analysis: Balancing Bandwidth and Computational Throughput
In the last section, we evaluated the three proposed dataflows and showed how data reuse enabled the OC optimization to significantly improve the runtime of HKS, saving considerable bandwidth even when streaming in $\mathbf{evks}$. In this section we conduct a sensitivity study to investigate how the performance of HKS with OC is affected as we increase both the bandwidth, ranging up to 1TB/s (HBM3), and the computational throughput by up to 16$\times$ more. We evaluate HKS performance for ARK and BTS3 benchmarks, the smallest and largest benchmarks, respectively. We refer to computational throughput as MODOPS (Modular Operations per Second).
@@ VI-C1 Bandwidth Analysis
In Figure 4(d) and (e) we increase the off-chip bandwidth beyond 64GB/s (DDR5) up to 1TB/s (HBM3) to study how scaling bandwidth affects HKS runtime for ARK and BTS3. For both ARK and BTS3, the benefit from OC compared to the other two dataflows diminishes at bandwidths larger than 256GB/s. At this point, the off-chip data movement is mostly masked by computation on the RPU, limiting further performance gain from increasing bandwidth.
According to Figure 4(e), moving from 1TB/s to the $OC_{base}$ BW for BTS3, results in a 31.25$\times$ bandwidth saving, with only 1.35$\times$ increase in runtime. On the other hand, transitioning to 32GB/s with the MP dataflow leads to a 13.98$\times$ slower runtime than the 1TB/s implementation.
In Figure 4(d), we see that employing OC dataflow for ARK, data movement becomes fully masked by computation after reaching a bandwidth of 128GB/s. In contrast, MP, operating at 1TB/s, still remains idle for 1.39$\times$ of its execution time. With 8GB/s of off-chip bandwidth and the OC implementation, we can save 16$\times$ bandwidth by being 1.6$\times$ slower compared to OC with 128GB/s. That is while for MP implementation, moving to 8GB/s causes a 5.17$\times$ increment in the runtime.
@@ VI-C2 Computational Throughput Analysis
Figure. 6: Evaluating HKS runtimes for ARK using OC at different MODOPS.
@@
The RPU design includes 128 lanes, meaning 128 modular multiplier, which is 128$\times$ less functional units compared to BTS [15]. To enhance the HKS runtime, which benefits from parallel computation where the calculation of each output tower is entirely independent of others, we simulate an accelerator with greater computational throughput by increasing RPU’s MODOPS (2$\times$, 4$\times$, 8$\times$, and 16$\times$). We will see how increasing the MODOPS will impact ARK’s OC performance and find the balance point between bandwidth and MODOPS. First, we will consider $\mathbf{evks}$ to be stored on-chip.
Based on Figure 6, for low bandwidths, the HKS runtime for different MODOPS are almost the same, as they are all bandwidth limited and do not benefit from increasing the computational throughput. However, with higher bandwidths, HKS starts to become compute bound, resulting in an increasing gap between the HKS runtime at different MODOPS. As mentioned earlier, in ARK’s OC implementation, off-chip data movement is entirely masked by computation at 128GB/s. We will call this point ARK’s saturation point. By doubling the MODOPS, we achieve similar performance with only 12.8GB/s bandwidth, saving 10$\times$ off-chip bandwidth. Table V outlines the required bandwidth and computational throughput for DC and MP to match ARK’s saturation point. To save bandwidth, both DC and MP require at least 4$\times$ more MODOPS. However, maintaining the MODOPS same as OC necessitates at least 4.26$\times$ and 10$\times$ higher bandwidth for DC and MP, respectively.
Figure 7 shows the required MODOPS and bandwidth to get the same performance as the baseline and ARK’s saturation point with OC while streaming the $\mathbf{evks}$ on-chip and having a 32MB on-chip memory. We previously showed that 12.8GB/s bandwidth with 2$\times$×MODOPS has the same performance as ARK’s saturation point, with $\mathbf{evks}$ being on-chip. To get the same performance while streaming the $\mathbf{evks}$, 2.6$\times$ more bandwidth with the same MODOPS is required, saving 16$\times$ on-chip memory. However, with only one RPU, 20$\times$ more bandwidth is required to get the saturation performance. At 256GB/s the design is completely compute bound; therefore, we can find a better ballance between bandwidth and MODOPS by increasing the MODOPS and decreasing the bandwidth. This intuition also applies to the baseline. As Figure 7(b) shows, by increasing the MODOPS we can achieve the same performance as 1$\times$×MODOPS and save 1.2$\times$ bandwidth.
@@ VII Related Work
High-Performance CPU Implementations: Many software libraries exist that support HE [7, 6, 8, 5]. Notably, Lattigo [7], OpenFHE [6], and HEAAN [5], support CKKS bootstrapping and FHE. However, due to limited computational power, some applications like deep neural network inference, remain impractical on pure CPU implementations. For instance, state-of-the-art ResNet models [18, 31, 32] can take upwards of one hour per inference, limiting their applicability to infrequent, low arrival rate applications [33, 34].
@@
Compilers and Dataflow: Recently, many works have focused on FHE compiler development, specifically in the context of private neural inference. Tools such as CHET [35] and EVA [36] automatically map and optimize common neural network layers for FHE and handle parameter selection and ciphertext management implicitly. Recently $\mathsf{Orion}$ [37] makes use of double-hoisting [38] to further improve CPU latencies. Recent work in MAD [25] highlights how clever caching techniques and dataflow optimizations can still yield high performance bootstrapping implementations despite small on-chip memories (1 to 32 MB). While MAD [25] presents a dataflow similar to our DC dataflow, we show our OC approach using 32 and 512MB of on-chip memory can further increase arithmetic intensity of key-switching by taking advantage of on-chip data reuse.
@@
GPU/FPGA Acceleration: GPUs offer promising opportunities for performance gain given their numerous parallel compute units and high off-chip bandwidth. Previous studies explore GPU acceleration of HE operations [39, 12, 40, 41, 42, 43]. Jung [41] was the first GPU implementation to support CKKS, suggesting optimizations like kernel fusing to reduce on-chip memory requirements. Still, the lack of native modular arithmetic support and limited on-chip memory yields subpar performance relative to modern ASIC solutions [13, 15, 16, 22].
@@
An alternative approach is to employ the use of FPGAs [44, 20, 45, 46]. While FPGAs may not provide the same level of performance as ASIC solutions, their reprogrammability offers a significant advantage given that FHE algorithms continue to evolve at a rapid pace. Poseidon [44] was the first FPGA-based accelerator targeting FHE CKKS. Most recently, FAB [20] proposes several dataflow optimizations, alongside an FPGA implementation, geared towards increasing on-chip data reuse and minimizing unnecessary DRAM accesses.
@@
ASIC Acceleration: Cheetah [47] was the first to present a large-scale ASIC with custom logic and show that the overheads of HE could be overcome with hardware acceleration. $\mathtt{F1}$ [10] presented a programmable ASIC accelerator for FHE, specifically tailored to small parameter sets with limited bootstrapping support, making it inefficient for deep neural network computations. Since then, many works [13, 15, 16, 22] proposed ASIC solutions targeting bootstrapping and FHE. CraterLake [13] and BTS [15] were the first ASIC accelerators supporting bootstrapping, but featured underutilized compute units despite allocating 1TB of off-chip bandwidth. Later, ARK [16] proposed a minimal key-switching strategy to address memory bottleneck, yet still required a large 512MB on-chip scratchpad. Recently, SHARP [22] characterized the effect of smaller moduli size on neural network accuracy, demonstrating the feasibility of using smaller 36-bit RNS moduli to reduce compute unit size, on-chip memory requirements, and off-chip bandwidth pressure.
@@ VIII Conclusion
In this paper, we propose three distinct approaches for implementing the hybrid key-switching algorithm that differ in the sequence of instructions and data reuse strategies. Through our evaluations using the RPU accelerator with various benchmarks, our novel dataflow, Output-Centric, achieves up to 4.16$\times$ speedup over a naive Max-Parallel implementation of HKS. Additionally, we can save 16$\times$ of on-chip SRAM by storing the $\mathbf{evks}$ off-chip. Here, OC achieves the same performance with 3.3$\times$ less bandwidth, and up to 2.4$\times$ more arithmetic intensity compared to MP with 64GB/s bandwidth and storing $\mathbf{evks}$ on-chip.

### v4

@@ VI-C Sensitivity Analysis: Balancing Bandwidth and Computational Throughput
In the last section, we evaluated the three proposed dataflows and showed how data reuse enabled the OC optimization to significantly improve the runtime of HKS, saving considerable bandwidth even when streaming $\mathbf{evks}$. In this section we conduct a sensitivity study to investigate how the performance of HKS with OC is affected as we increase both the bandwidth, ranging up to 1TB/s (HBM3), and the computational throughput by up to 16$\times$ more. We evaluate HKS performance for ARK and BTS3 benchmarks, the smallest and largest benchmarks, respectively. We refer to computational throughput as MODOPS (Modular Operations per Second).
@@ VI-C1 Bandwidth Analysis
In Figure 4(d) and (e) we increase the off-chip bandwidth beyond 64GB/s (DDR5) up to 1TB/s (HBM3), with HPLEs fixed at 128, to study how scaling bandwidth affects HKS runtime for ARK and BTS3. For both ARK and BTS3, the benefit from OC compared to the other two dataflows diminishes at bandwidths larger than 256GB/s. At this point, the off-chip data movement is mostly masked by computation on the RPU, limiting further performance gain from increasing bandwidth.
According to Figure 4(e), moving from 1TB/s to the $OC_{base}$ BW for BTS3, results in a 31.25$\times$ bandwidth saving, with only a 1.35$\times$ increase in runtime, while transitioning to the same bandwidth (32GB/s) with the MP dataflow leads to a 13.98$\times$ slower runtime than the 1TB/s implementation.
In Figure 4(d), we see that by employing OC dataflow for ARK, data movement becomes fully masked by computation after reaching a bandwidth of 128GB/s. With 8GB/s of off-chip bandwidth ($OC_{base}$) and the OC implementation, we can save 16$\times$ bandwidth by being 1.6$\times$ slower compared to OC with 128GB/s. That is while for MP implementation, moving to 8GB/s causes a 5.17$\times$ increment in the runtime.
@@ VI-C2 Computational Throughput Analysis
Figure. 8: Evaluating HKS runtimes for ARK using OC at different MODOPS, with evks on-chip.
@@
The RPU design includes 128 lanes, meaning 128 modular multipliers, which is 128$\times$ less functional units compared to BTS [16]. To enhance the HKS runtime, which benefits from parallel computation where the calculation of each output tower is entirely independent of others, we simulate an accelerator with greater computational throughput by increasing the RPU’s MODOPS (2$\times$, 4$\times$, 8$\times$, and 16$\times$). We will see how increasing the MODOPS will impact ARK’s OC performance. First, we will consider $\mathbf{evks}$ to be pre-loaded on-chip.
Based on Figure 8, at low bandwidths, the HKS runtime for different MODOPS is nearly identical, as they are bandwidth limited and do not benefit from increased computational throughput. However, at higher bandwidths, HKS becomes compute bound, resulting in an increasing gap between the HKS runtime at different MODOPS. As mentioned earlier, in ARK’s OC implementation, off-chip data movement is entirely masked by computation at 128GB/s. We will call this point “ARK’s saturation point”. At 128GB/s, the design is no longer limited by bandwidth. Doubling the MODOPS reduces the HKS runtime. Therefore, as shown in Figure 8, saturation performance can be achieved with less bandwidth (12.8GB/s) with 2$\times$ MODOPS, saving 10$\times$ on-chip bandwidth.
Table V outlines the required bandwidth and computational throughput for DC and MP to match ARK’s saturation point. Maintaining the same MODOPS as OC necessitates at least 4.26$\times$ and 10$\times$ higher bandwidth for DC and MP, respectively.
Figure 9 shows the required MODOPS and bandwidth to get the same performance as “ARK’s saturation point” and the “baseline” with OC while streaming the $\mathbf{evks}$ on-chip and having a 32MB on-chip memory. We previously showed that 12.8GB/s bandwidth with 2$\times$×MODOPS has the same performance as ARK’s saturation point, with $\mathbf{evks}$ being on-chip. To get the same performance while streaming the $\mathbf{evks}$, 2.6$\times$ more bandwidth with the same MODOPS (2$\times$) is required, saving 12.25$\times$ on-chip memory. However, with 1$\times$×MODOPS, 20$\times$ more bandwidth is required to get the saturation performance. At 256GB/s the design is completely compute bound; therefore, we can find a better balance between bandwidth and MODOPS by increasing the MODOPS and decreasing the bandwidth. This intuition also applies to the baseline. As Figure 9(b) shows, by doubling the MODOPS we can achieve the same performance as 1$\times$×MODOPS and save 1.2$\times$ bandwidth.
@@ VII Related Work
High-Performance CPU Implementations: Many software libraries exist that support HE [7, 6, 8, 5]. Notably, Lattigo [7], OpenFHE [6], and HEAAN [5] support CKKS bootstrapping and FHE. However, due to limited computational power, some applications like deep neural network inference, remain impractical on pure CPU implementations. For instance, state-of-the-art ResNet models [19, 31, 32] can take upwards of one hour per inference, limiting their applicability to infrequent, low arrival rate applications [33, 34].
@@
Compilers and Dataflow: Recently, many works have focused on FHE compiler development [35, 36, 37, 38, 26, 39, 40, 41, 42], particularly for private neural inference. Tools such as CHET [35] and EVA [36] automatically map and optimize common neural network layers for FHE, managing parameter selection and ciphertexts implicitly. HECO [37] is an end-to-end compiler that converts high-level programs into secure FHE circuits, allowing non-experts to develop secure and efficient FHE applications. Porcupine [39] and Coyote [42] generate vectorized HE code for small HE kernels, with Coyote [42] optimizing data layout to minimize the number of rotations. $\mathsf{Orion}$ [38] utilizes double-hoisting [43] to enhance CPU latencies. MAD [26] highlights how caching techniques and dataflow optimizations can yield high performance bootstrapping implementations despite small on-chip memories (1 to 32 MB). While MAD [26] presents a dataflow similar to our DC approach, our OC strategy, using 32 and 392MB of on-chip memory, further increases the arithmetic intensity of key-switching by leveraging on-chip data reuse.
@@
GPU/FPGA Acceleration: GPUs offer performance gains with their numerous parallel compute units and high off-chip bandwidth. Previous studies explore GPU acceleration of HE operations [44, 12, 45, 46, 47, 48]. Jung [46] was the first the support CKKS on GPU, suggesting optimizations like kernel fusing to reduce on-chip memory requirements. Still, the lack of native modular arithmetic support and limited on-chip memory yields subpar performance compared to modern ASIC solutions [14, 16, 17, 23].
@@
Another approach is to use FPGAs [13, 49, 21, 50, 51]. While FPGAs may not provide the same level of performance as ASIC solutions, their re-programmability is advantageous as FHE algorithms evolve. HEAX [13] is an FPGA-based accelerator for FHE CKKS. Recently, FAB [21] and Poseidon [49] proposed FPGA-based accelerators to support bootstrapping. FAB [21] proposes several dataflow optimizations, alongside an FPGA implementation, geared towards increasing on-chip data reuse and minimizing unnecessary DRAM accesses.
@@
ASIC Acceleration: Cheetah [9] was the first to present a large-scale ASIC with custom logic and show that the overheads of HE could be overcome with hardware acceleration. $\mathtt{F1}$ [10] presented a programmable ASIC accelerator for FHE but was tailored to small parameter sets with limited bootstrapping support, making it inefficient for deep neural network computations. Since then, many works [14, 16, 17, 23] proposed ASIC solutions targeting bootstrapping and FHE. CraterLake [14] and BTS [16] were the first ASIC accelerators supporting bootstrapping, but featured underutilized compute units despite allocating 1TB of off-chip bandwidth. Later, ARK [17] proposed a minimal key-switching strategy to address memory bottlenecks, yet still required a large 512MB on-chip scratchpad. Recently, SHARP [23] characterized the effect of smaller moduli size on neural network accuracy, demonstrating the feasibility of using smaller 36-bit RNS moduli to reduce compute unit size, on-chip memory requirements, and off-chip bandwidth pressure.
@@ VIII Conclusion
In this paper, we propose three distinct approaches for implementing the hybrid key-switching (HKS) algorithm that differ in the sequence of instructions and data reuse strategies. The large amount of $\mathbf{evks}$ and generated intermediate data (up to 1.5GB) puts pressure on implementing FHE applications. We introduce a novel dataflow, named Output-Centric (OC), that reduces the off-chip data movement and increases data reuse. Through our evaluations using the RPU accelerator with various benchmarks, our OC dataflow achieves up to 4.16$\times$ speedup over a naive Max-Parallel (MP) implementation of HKS. Additionally, we can save 12.25$\times$ of on-chip SRAM by storing the $\mathbf{evks}$ off-chip and save 3.3$\times$ bandwidth with up to 2.4$\times$ more arithmetic intensity compared to a MP on-chip implementation.
Acknowledgment
Support for this work was provided in part by NSF CAREER award #2340137. This work was supported in part by Graduate Assistance in Areas of National Need (GAANN) and funding from DARPA, under the Data Protection in Virtual Environments (DPRIVE) program, contract HR0011-21-9-0003. Reagen and Ebel received generous support from the NY State Center for Advanced Technology in Telecommunications (CATT) and a gift award from Google. We also express our gratitude to Duality, especially David Bruce Cousins, Ahmad Al Badawi, and Yuriy Polyakov as well as Andrew G. Schmidt from AMD for their valuable insights and support. The views, opinions, and/or findings expressed are those of the authors and do not necessarily reflect the views of sponsors.